E15-3 Mama Leone's Frozen Pizzas uses 50,000 units of cheese per year. Each unit costs $2.50. The ordering cost for the cheese is $250 per order, and its carrying cost is $0.50 per unit per year. Calculate the firm's economic order quantity (EOQ) for the cheese. Mama Leone's operates 250 days per year and maintains a minimum inventory level of 2 days' worth of cheese as a safety stock. If the lead time to receive orders of cheese is 3 days, calculate the reorder point.
Answer #1

Economic order quantity = (2 x 250 x 50000/0.5)1/2 = 7071

Usage per day = 50000/250 = 200 units of cheese

Reorder point = 3 days x 200/day + 2 days x 200/day = 1000 units

Hope this helps!
Answer #2

EOQ =√ 2 x annual demand x order cost = √2 x 125000 x 250 = 50units

                 √Carrying cost                            √25000

Reorder point = lead time demand + safety stock= 3 + 2 = 5 days
Answer #3

Mama Leone's Frozen Pizzas uses 50,000 units of cheese per year. Each unit costs $2.50. The ordering cost for the cheese is $250 per order, and its carrying cost is $0.50 per unit per year.

Calculate the firm's economic order quantity (EOQ) for the cheese.

(2*50000*250/.5)^.5 = 7071 units is the EOQ.

========================================

Mama Leone's operates 250 days per year and maintains a minimum inventory level of 2 days' worth of cheese as a safety stock. If the lead time to receive orders of cheese is 3 days, calculate the reorder point.

Consumption per day = 50000/250 = 200

Reorder point = 5*200 = 1000 units
Answer #4
Mama Leone’s Frozen Pizzas uses 50,000 units of cheese per year. Each unit costs $2.50. The ordering cost for the cheese is $250 per order, and its carrying cost is $0.50 per unit per year. Calculate the firm’s economic order quantity (EOQ) for the cheese. Mama Leone’s operates 250 days per year and maintains a minimum inventory level of 2 days’ worth of cheese as a safety stock. If the lead time to receive orders of cheese is 3 days, calculate the reorder point.
answered by: Hussain